OpenStudy (anonymous):

how do i find the vertical asymptote of the function f(x)=4/(x+5)

OpenStudy (amistre64):

you first cross out like terms from top to bottom, then what ever value makes the bottom a zero is your VA

OpenStudy (amistre64):

in this case....x =-5

OpenStudy (anonymous):

so you just make x+5=0

OpenStudy (anonymous):

then solve for x?

OpenStudy (anonymous):

Try finding the lim as x goes to 0

OpenStudy (amistre64):

that is correct, you want to know what vale of x can never be used

OpenStudy (amistre64):

x=-5 is a nono in this equation, so never use it, draw a lint to mark its position and stay away from it, that your VA

OpenStudy (amistre64):

line not lint, but if it is in macaroni art, use zitis

OpenStudy (anonymous):

so to find the equation of the vertical asymptote of the function f(x)=4/x+5 what is the answer?

OpenStudy (anonymous):

i am so confused?

OpenStudy (amistre64):

lie down on my couch right here and lets talk about this confusion :) tell me, can the bottom of a fraction ever be a zero?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

then the line on a graph that marks the spot where the bottom would be zero is indictaed by the vertical asymptote. at the value of x that is bad.

OpenStudy (amistre64):

what value of x makes the bottom of your fraction there go bad?

OpenStudy (anonymous):

anything negative?

OpenStudy (amistre64):

negative numbers are not bad numbers they are just misunderstood :) there is on value in particular that is going to be offlimits: x+5 = 0 when x =?

OpenStudy (anonymous):

-5

OpenStudy (amistre64):

correct! so that is your VA

OpenStudy (amistre64):

x = -5 is the VA

OpenStudy (anonymous):

ok i get it now,

OpenStudy (anonymous):

there i graphed in it matlab shows you what is happening in case the visual aspect helps

OpenStudy (amistre64):

you got that in crayon? :)

OpenStudy (anonymous):

then how do i get horizontal asym to 5x^2-4/(x+1)

OpenStudy (amistre64):

we divide the top by the bottom like noremal long division to get an exact answer, but its not a horizontal asymptote.... itll be a line that is slanted

OpenStudy (anonymous):

look at the graph that should help you visualize it.

OpenStudy (amistre64):

the right side there seems off, but that might be the zoom limit

OpenStudy (anonymous):

ill try to fix it

OpenStudy (amistre64):

if we simply do a shortcut and divide all the numbers by x, we get the top to be 5x

OpenStudy (anonymous):

there you go now its a bit better

OpenStudy (amistre64):

thats better, but the your asymptotes gone parabolic instead of linear :)

OpenStudy (amistre64):

5x^2-4/(x+1) the top and bottom differ by 1 degree, so the asymptote is linear

OpenStudy (anonymous):

sorry i had the equation not copy and pasted wrong lol here is the correct result

OpenStudy (amistre64):

yay!! i knew you could do it :)

OpenStudy (anonymous):

man im just trying to help him out?

OpenStudy (amistre64):

youre doing great....

OpenStudy (anonymous):

im pretty sure when you learn these things it helps to look at a graph

OpenStudy (amistre64):

it does, i am visual alot, the analysis part still gets my brain tied in knots

OpenStudy (anonymous):

sorry i thought you were being sarcastic there

OpenStudy (amistre64):

;) wel it was in humour, but nothing cruel or hateful. we all have errors to keep us humble i beleive, i know i have my share

OpenStudy (anonymous):

well to be honest i did have the right equation i just got excited using matlab and didn bracket off the exponents properly